# RadMamba: Efficient Human Activity Recognition through Radar-based Micro-Doppler-Oriented Mamba State-Space Model

### v1

@@ IV-C Solution 3: Convolution Projections in rMamba Blocks
$\mathbf{X}_{bw}$ $=\text{flip state dim}(\mathbf{X}_{fw}),$ (12)
where $\mathbf{Z},\mathbf{X}_{fw},\mathbf{X}_{bw}\in\mathbb{R}^{N\times dim}$. Both Conv1D operations utilize a kernel size (in channels, out channels, kernel size) = $(dim,dim,3)$.
When $\mathbf{X}_{fw}$ and $\mathbf{X}_{bw}$ pass through the SSM branches with architecture of:
[…]
The outputs from both directions are combined using a gating mechanism modulated by the SiLU-activated $\mathbf{Z}$ and summed:
$\mathbf{Y^{\prime}}_{fw/bw}$ $=\mathbf{Y}_{fw/bw}\odot\text{SiLU}(\mathbf{Z})$ (14)
$\mathbf{Y}_{sum}$ $=\mathbf{Y^{\prime}}_{fw}+\mathbf{Y^{\prime}}_{bw}$ (15)
where $\odot$ denotes element-wise multiplication. Additionally, in our proposed rMamba, the projection layer P3 is also replaced by Conv1D, so that:
$\mathbf{Y}_{out}=\text{Conv1D}_{P3}(\mathbf{Y}_{sum})+\mathbf{X}_{proj},$ (16)
where the final Conv1D has weights of dimensions $(dim,dim,1)$ and transforms the combined sequence back to $(N,dim)$.
These convolutional layers enable the model to capture subtle variations in Doppler-span changes over time, which are crucial to describe the kinematic evolution of the different activities to be classified. Also, let us take Dataset CI4R as a comparison reference and follow the method to calculate cross-correlation as explained in Section III Issue 3. While passing through RadMamba, the average time bin cross-correlation is 0.002 before projection 1 and 2.
Then, compared to conventional linear projection, the convolution projection 1 increased the averaged cross correlation gap between projection input and output from 77.9 to 1065.7. The convolutional projection 2 increased this gap from 0.4 to 16.1. The convolutional projection 3 increased the average time cross-correlation from 245.3 to 356.4, while the linear projection 3 decreased this number. This enhancement significantly improves the model’s sensitivity to Doppler patterns between patches and makes it easier to regard all patches as one activity, thereby boosting classification performance.
IV-D Selective State-Space Modeling for Micro-Doppler Sequences
Figure 3: The architecture of State Space Model in RadMamba.
The SSM is a significant layer in rMamba enables efficient sequence modeling for the dynamic nature of micro-Doppler signatures. As shown in Fig. 3 (a), the discrete SSM is defined as:
$\mathbf{y}[n]$ $=\mathbf{C}\mathbf{h}[n]+\mathbf{D}\mathbf{x}[n],$ (18)
$\bar{\mathbf{A}}$ $=\exp(\boldsymbol{\Delta}\cdot\mathbf{A}),$ (19)
[…]
The input of the SSM branches in the proposed rMamba block are tensors $\mathbf{X}_{fw/bw}\in\mathbb{R}^{N\times dim}$; after Conv1D and LayerNorm in SSM branch, this is termed as $\mathbf{X^{\prime}}_{fw/bw}\in\mathbb{R}^{N\times dim}$.
Therefore, the dimension of the discretized state transition ${\mathbf{\bar{A}}}$ is ${N\times dim\times{dim_{s}}}$ and the input matrices dimension ${\mathbf{\bar{B}}}$ is ${N\times N\times{dim_{s}}}$, respectively. While $N$ equals the time dimension in micro-Doppler signatures, the input matrices $\mathbf{\bar{B}}$ captures how the Doppler-aligned vector influences the state $\mathbf{h}[n]$, and the discretized state transition weight ${\mathbf{\bar{A}}}$ capture how the last state $\mathbf{h}[n-1]$. Moreover, the matrices $\mathbf{C}\in\mathbb{R}^{N\times dim_{s}}$ translate the current state $\mathbf{h}[n]$ to the output. The skip matrices $\mathbf{D}\in\mathbb{R}^{dim}$ add weighted input directly to output of SSM, which helps the recurrent part focus on the temporal evolution.
$\mathbf{B}$ $=\mathbf{W}_{\mathbf{B}}\mathbf{X},$ (21)
$\mathbf{C}$ $=\mathbf{W}_{\mathbf{C}}\mathbf{X},$ (22)
[…]
IV-E Overall Architecture
To summarize the method, the proposed RadMamba pipeline in pseudo-code is presented in Algorithm I. The patch embedding in PEmb follows the method in  [15] and the position embedding method is a sinusoidal position encoding. [32].
Algorithm 1 Proposed RadMamba Pipeline
Parameters: $dim$, $dim_{s}$, $dt\_rank$
1:Channel Fusion and Downsample:
2:   $\mathbf{X_{c-d}}=\textit{Chan-DS}(\mathbf{X})$, $\mathbf{X}_{c-d}\in\mathbb{R}^{C_{c-d}\times H_{c-d}\times W_{c-d}}$
3:Doppler-Aligned Segmentation:
4:   $\mathbf{X_{seg}}=\textit{DA-Seg}(\mathbf{X}_{c-d})$, $\mathbf{X}_{seg}\in\mathbb{R}^{C_{c-d}\times N\times H_{c-d}}$
5:Patch and Position Embedding:
6:   $\mathbf{X^{\prime}}_{p}=\textit{PEmb}(\mathbf{X}_{seg})$, $\mathbf{X^{\prime}}_{p}\in\mathbb{R}^{N\times D}$
7:rMamba Block:
8:   $\mathbf{Y}_{out}=\textit{rMamba}(\mathbf{X}_{p})$, $\mathbf{Y}_{out}\in\mathbb{R}^{N\times dim}$
9:Classify:
10:   $\mathbf{F}=\mathbf{W}_{out}\mathbf{Y}_{out}$, $\mathbf{F}\in\mathbb{R}^{Q}$
V Open-source Datasets and Preprocessing for Radar-based HAR
To evaluate the efficacy of our proposed RadMamba framework, we utilize three distinct radar datasets. This section describes these datasets, namely the CW non-continuous Dataset DIAT, FMCW non-continuous Dataset CI4R, and FMCW continuous Dataset UoG2020, along with their respective pre-processing steps. More details on these datasets are summarized in Tab. I. Here the word ‘continuous’ refers to samples of data where sequences of human activities are performed one after the other in a continuous manner.
TABLE I: Information of Different Radar-based Activity Recognition Datasets Used in This Work.
Datasets DIAT CI4R UoG2020
Continuous No No Yes
Classes 6 11 6
Sensor X-band CW IWR1443 FMCW FMCW
Freq.(GHz) 10 77 5.8
Resolution 224$\times$224 224$\times$224 240$\times$1471
BW(MHz) - 750 400
Length per sample (s) 3 20 35
Train Samples 2646 528 39
Test Samples 759 132 6
Sliding Window 224
V-A Dataset DIAT: X-Band CW Radar with Discrete Activity Instances [30]
This dataset was collected using an X-band CW radar operating at a frequency of 10 GHz. It encompasses data from 30 human subjects performing ‘suspicious’ activities at ranges spanning 10 meters to $\pm$0.5 kilometers over 3-second durations, resulting in an imbalanced set of 3,780 spectrogram samples of shape $(3,244,244)$across six activity classes. Preprocessing adheres to the methodology outlined in [30]. To facilitate a fair comparison with previous works, the dataset is partitioned into 70% training, 10% validation, and 20% test sets within each class, maintaining class-specific balance for reliable performance evaluation.
V-B Dataset CI4R: Short-Range FMCW Radar for Multi-Class Activity Recognition [31]
The experiments of the second dataset employ a Texas Instruments IWR1443 FMCW radar, operating at 77 GHz with a bandwidth of 750 MHz. It includes data from six participants, varying in age, height, and weight, performing 11 distinct activities and ambulatory gaits. Each participant executed 10 repetitions per activity, yielding 60 samples per class per sensor, as detailed in [31]. Preprocessing involves applying a 256-point short-time Fourier transform (STFT) with a context window length to convert the raw radar signals into a 2D micro-Doppler representation of shape $(1,244,244)$. The dataset is split with an 8:2 training-to-test ratio, ensuring sufficient data for model optimization and evaluation.
V-C Dataset UoG2020: Continuous FMCW Radar for Realistic HAR Sequences [27]
Representing a leap in complexity, this dataset was acquired using an FMCW radar operating at 5.8 GHz with a 400 MHz bandwidth. It comprises data from 15 participants (14 males, 1 female, aged 21–35) performing six basic activities within continuous 35-second sequences. These activities are organized into three distinct sequential orders, with participants granted autonomy in selecting transition points between consecutive activities and determining trajectories for translational tasks, as described in [27]. The resulting micro-Doppler signatures are preprocessed into $(1,240,244)$ 2D frames, with two participants’ data designated as the test set, while the remaining 13 contribute to training, providing a challenging benchmark for continuous sequence modeling.
VI Experimental Setup
VI-A Training Hyperparameter
Training is conducted using the ADAMW optimizer [33, 34], paired with the ReduceLROnPlateau scheduler to adjust the learning rate based on test performance adaptively. For Dataset DIAT, the initial learning rate is set to $1\times 10^{-4}$, with a batch size of 16, and training proceeds for 50 epochs. For Dataset CI4R, RadMamba employs an initial learning rate of $5\times 10^{-3}$, while the baseline models use $1\times 10^{-4}$, with a batch size of 16 over 100 epochs, reflecting the dataset’s increased complexity and sample size. For Dataset UoG2020, the initial learning rate is $5\times 10^{-5}$, with a larger batch size of 256 and 50 epochs, processing continuous sequences with a frame length of 224 and a stride of 1 to capture temporal dependencies effectively. All experiments are executed on a single NVIDIA RTX 4090 GPU.
To account for training variability, all models, including RadMamba, are trained and evaluated ten times with distinct seeds, reporting in the following analysis the mean and standard deviation of performance metrics in the model size swept experiment.
Figure 4: Accuracy vs. number of parameters (1) and number of ##\##FLOP per inference (2) on: (a) Dataset DIAT (b) Dataset CI4R (c) Dataset UoG2020
VI-B Network Structure
TABLE II: Network Architecture of Models Used in This Work
Model Swept Hidden Size ($dim$ for RadMamba)
ResNet [5, 7, 12, 20, 30]
Bi-LSTM [1, 2, 4, 10, 20, 28, 34]
CNN-LSTM [2, 4, 6, 16, 28, 34, 46]
CNN-Bi-GRU [1, 2, 4, 6, 16, 28]
RadMamba Dataset DIAT [8, 16, 32, 64, 80]
RadMamba Dataset CI4R [8, 16, 32, 64, 80, 96, 128, 160]
RadMamba Dataset UoG2020 [8, 16, 20, 24, 32]
TABLE III: Specified Network Architecture Variations of RadMamba
Datasets $dim_{s}$ $dt\_rank$ Chan. Fusion DS Redu. Factor
Dataset DIAT 1 2 [2, (16, 3, 3)] (2, 2)
Dataset CI4R 4 0 [1, (1, 3, 3)] (2, 8)
Dataset UoG2020 16 4 [1, (1, 3, 3)] (2, 32)
The classification performance of RadMamba is evaluated against a suite of established models, selected for their proven efficacy across CNN, RNN, and hybrid architectures: ResNet [35], Bi-LSTM [27], CNN-LSTM [13], and CNN-Bi-GRU [14]. To explore the impact of model scaling, we conducted a model size swept experiment, varying the hidden dimensions as shown in Tab. III. These sizes are chosen to capture the full spectrum of performance trends as model capacity increases, providing insights into scalability and efficiency.
Notably, the CNN-Bi-GRU model, originally designed for distributed radar systems in  [14], exhibits a larger parameter count, though its channel fusion approach may not generalize across all datasets.
For RadMamba, specific configurations of the channel fusion and down-sampling blocks, as well as the SSM parameters, are tailored to each dataset to optimize micro-Doppler processing as shown in Tab. III, where the ‘channel fusion’ column specifies [L, ($C_{c-d},H_{k},W_{k}$)], where L is the number of channel fusion blocks and the $C_{c-d},H_{k},W_{k}$ are channel, kernel height, kernel width of Conv2D layer, respectively. The ‘DS Redu. Factor’ denotes the reduction factors ($H/H_{c-d}$, $W/W_{c-d}$).
For transformer-based and SSM-based models such as FML-ViT [16], LW-ViT [21], and ActivityMamba [19], we adopt their reported results as optimal outcomes under minimal model sizes, as dimension definitions may vary and are not always fully specified in the original works.
VII Results and Discussion
VII-A Model Scaling and Efficiency Analysis
The model size swept results, illustrated in Figure 4, compare accuracy against the number of parameters from 2k to 150k and ##\##FLOP per inference across Datasets DIAT, CI4R, and UoG2020 between previous models ResNet, Bi-LSTM, CNN-Bi-GRU, and CNN-LSTM and the proposed RadMamba.
For Dataset DIAT, as presented in Fig. 4 (a1) and (a2), RadMamba achieves a peak accuracy of 99.8% with only 21.7 k parameters. However, RadMamba overfits with a model size larger than 21.7 k because the dataset is very simple and the accuracy is close to 100%. In contrast, ResNet’s accuracy increases steadily with model size, reaching 99.9% with 208.1 k parameters, yet it requires nearly ten times the parameters of RadMamba for slightly lower performance.
TABLE IV: Mean and Standard Deviation of Classification Accuracy Performance across Seed 0 to 9 of Different NN-based RadHAR Models Evaluated With Three Different Radar Settings alongside Their Model Size and Floating-point Operations per Inference Sample (FLOP/Inf.)
Dataset DIAT [30]
(CW, Non-continuous)
Dataset CI4R [31]
(FMCW, Non-continuous)
Dataset UoG2020 [27]
(FMCW, Continuous)
Classifiers
#params
#FLOP/Inf. a
Accuracy
#params
#FLOP/Inf. a
Accuracy
#params
#FLOP/Inf. a
Accuracy
ResNetb [35] 69.4 253.3 99.8$\pm$0.16 150.7 310.2 32.3$\pm$3.35 24.7 72.8 83.6$\pm$1.73
Bi-LSTM [27] 23.3 5.6 93.9$\pm$0.68 134.4 6.0 79.4$\pm$1.23 15.5 2.1 85.3$\pm$2.01
CNN-LSTM [13] 115.3 8.7 98.4$\pm$0.33 96.5 8.7 86.1$\pm$1.32 10.3 1.2 86.1$\pm$2.22
CNN-Bi-GRU [14] 55.0 1090.1 98.5$\pm$0.41 142.8 1013.0 84.9$\pm$1.11 53.6 1082.0 85.4$\pm$2.19
FML-ViTc [16] - - - $\sim$2700 169.0 $\sim$92.0 - - -
LW-VITc [21] - - - 769 $\sim$2410 $\sim$92.0 - - -
ActivityMambac [19] $\sim$8700.0 $\sim$1220.0 99.8 - - - - -
RadMamba (Ours) 21.7 145.6 99.8$\pm$0.12 71.4 8.8 91.2$\pm$2.57 6.7 1.0 89.3$\pm$1.38
RadMamba (Ours) 10.1 14.5 98.2$\pm$0.46 47.9 6.0 88.8$\pm$4.58 17.7 1.8 89.6$\pm$0.78
@@ VII-A Model Scaling and Efficiency Analysis
    ##\##FLOP/Inf. is the number of Floating point operations per inference.
@@
    This ResNet only retains the structure of ResNet18 from the first layer to the end of the second residual block. The relative multiples of channels between convolution layers remain unchanged, but the size of the minimum channel can be adjusted.
@@
Figure 5: Comparison between ground truth and best prediction of RadMamba model with the configuration in Table I on (a) Dataset DIAT; (b) Dataset CI4R; (c) Dataset UoG2020.
On Dataset CI4R, as shown in Fig. 4 (b1) and (b2), RadMamba consistently outperforms previous models across a range of parameter configurations in efficiency, maintaining superior accuracy with fewer parameters and ##\##FLOP/Inf. Similarly, for Dataset UoG2020 shown in Fig. 4 (c1) and (c2), RadMamba still records a higher accuracy for continuous activities scenarios with fewer ##\##FLOP/Inf., outperforming larger and more computationally intensive models in efficiency and effectiveness.
These trends, as depicted in the figure, affirm RadMamba’s capability to achieve superior accuracy across diverse radar modalities while drastically reducing both parameter count and ##\##FLOP/Inf., positioning it as an optimal solution for lightweight, real-time human activity recognition.
VII-B Comparison with Previous Works
Figure 6: Averaged test loss across 10 random seeds vs. Epochs during training of models with configuration in Table I on (a) Dataset DIAT; (b) Dataset CI4R; (c) Dataset UoG2020.
@@ VII-B Comparison with Previous Works
The experimental results detailed in Table IV highlight the superior performance of our proposed RadMamba model across multiple radar-based human activity recognition datasets.
On Dataset DIAT, which utilizes CW radar in a non-continuous setting, RadMamba achieved an accuracy of 99.8% with a standard deviation of $\pm$0.12%, using only 21.7k parameters and requiring 145.6 million ##\##FLOP/Inf. per sample. This demonstrates its capability to match the state-of-the-art accuracy of previous models while being significantly more parameter-efficient compared to $\sim$8700k parameters and computationally lighter compared to $\sim$1220M ##\##FLOP per inference. The CNN-LSTM with 8.7 M ##\##FLOP per inference achieves better accuracy of 98.4% than 98.2% from RadMamba with 14.5 M ##\##FLOP per inference. However, take 45nm technology for example, the energy consumption for 8kB SRAM memory access of one 32-bit floating point parameter and one 32-bit floating point computing operation is 5.0 pJ and 4.6 pJ [36]. With 11$\times$ larger model and around half ##\##FLOP per inference, it’s hard for CNN-LSTM to be more energy-efficient than RadMamba.
@@
TABLE V: Ablation Study with progressive techniques (denoted by Tech.) of New Blocks in RadMamba across Seed 0 to 5, Evaluated with Three Different Radar Settings alongside Their Number of Parameters
Dataset DIAT Dataset CI4R Dataset UoG2020
@@
    $(H_{1},W_{1})$ equals to (7, 7) for Dataset DIAT and B, (5, 7) for Dataset UoG2020 due to the different image size of these datasets.
@@
    The kernel size for projection 1 and 2 is $(dim,3,3)$, and for projection 3 is $(dim,1,1)$.
For Dataset CI4R, employing FMCW radar in a non-continuous scenario with 11 activities, RadMamba attained an accuracy of 91.2% with a standard deviation of $\pm$2.57%, leveraging 71.4 thousand parameters. This performance outperforms previous models shown in Tab. IV. Moreover, it remains competitive with the state-of-the-art of previous models while requiring substantially fewer computational resources (8.8 M ##\##FLOP/Inf. vs. 169 M to 2410 M ##\##FLOP/Inf.), showcasing its ability to excel in more complex activity recognition tasks with minimal computational overhead. Moreover, compared to the Bi-LSTM model with lowest arithmetic intensity in table, RadMamba outperforms Bi-LSTM accuracy of 9.4% with 86.5 k less parameters and same ##\##FLOP per inference of 6.0 M.
On Dataset UoG2020, which involves FMCW radar in a continuous setting with 6 activities, RadMamba recorded an accuracy of 89.3% with a standard deviation of $\pm$1.38%, utilizing 6.7 thousand parameters and 1.3M ##\##FLOP per inference. With an even smaller model size and more complicated task compared to Datasets DIAT and CI4R, RadMamba achieves the best accuracy in models shown in Tab. IV and highlights its effectiveness in continuous HAR scenarios with the smallest model size and ##\##FLOP per inference.
Figure 5 (a) and (b) present confusion matrices comparing the ground truth labels with the best predictions of RadMamba, illustrating its classification accuracy and error distribution for each dataset. Figure 5 (c) shows the comparison between the classification results of RadMamba and CNN-LSTM over time. Complementing this, Figure 6 displays the averaged test loss across ten random seeds as a function of training epochs, providing insight into the model’s convergence behavior and stability during training on the same datasets.
VII-C Ablation study of the Micro-doppler-oriented Structure
@@
Table V presents an ablation study evaluating the impact of new architectural blocks in the RadMamba model across three datasets with varying radar settings. The study systematically assesses the effects of different projection layers, patch sizes, and down-sampling scales on model accuracy and parameter efficiency, with results averaged over seeds 0 to 5.
@@
At the outset, the most important conclusion is that optimal performance emerges from the combined integration of all three proposed new blocks—projection layer, patch size, and down-sampling scale—rather than any single component in isolation. This is evident in the incremental improvements observed across the configurations. For instance, the baseline conventional ViM model (Row 1) achieves an accuracy of 93.5% on Dataset DIAT, 61.9% on Dataset CI4R, and 71.2% on Dataset UoG2020. In contrast, the proposed RadMamba (Row 27), which incorporates a 1D convolution layer, a patch size of ($H_{c-d}$, 1), and a down-sampling (2, 2), achieves a higher accuracy of 99.8% on Dataset DIAT. Similarly, on Dataset CI4R, Row 27 with a (2, 8) down-sampling scale yields significantly higher accuracy of 89.6%, and on Dataset UoG2020 with a (2, 32) scale, it reaches 88.6%. Intermediate configurations, such as Row 19 (only projection replacement), Row 7 (only patch size changing) and Row 2 (only down-sampling applied) show lower accuracies—underscoring that no single structure alone achieves the peak performance of the fully combined design in Row 27.
@@
To provide a more detailed analysis, we discuss the three blocks—projection layer, patch size, and down-sampling scale—separately, evaluating their individual contributions to RadMamba’s performance across the datasets.
VII-C1 Projection Replacement

### v3

@@ VI-A Comparison with Previous Works
    #\#FLOP. is the number of floating-point operations per inference.
@@
    This ResNet6 only retains the first 5 layers and the output linear layers of ResNet18.
@@
Figure 5: Example of comparisons between ground truth and best prediction of RadMamba model with the configuration in Table I on (a) Dataset DIAT; (b) Dataset CI4R; (c) Dataset UoG20.
@@
The experimental results detailed in Table V highlight the superior performance of our proposed RadMamba model across multiple radar-based human activity recognition datasets.
Figure 5 (a) and (b) present confusion matrices comparing the ground truth labels with the best predictions of RadMamba, illustrating its classification accuracy and error distribution for each dataset. Figure 5 (c) shows the comparison between the classification results of RadMamba and CNN-LSTM over time.
[…]
For Dataset CI4R, employing FMCW radar in a non-continuous scenario with 11 activities, RadMamba attained an accuracy of 91.2% with a standard deviation of $\pm$2.57%, leveraging 71.4 thousand parameters. This performance outperforms previous models shown in Tab. V. Moreover, it remains competitive with the state-of-the-art of previous models while requiring substantially fewer computational resources (8.8 M #\#FLOP/Inf. vs. 169 M to 2410 M #\#FLOP/Inf., showcasing its ability to excel in more complex activity recognition tasks with minimal computational overhead. Moreover, compared to the Bi-LSTM model, which has the lowest arithmetic intensity among the baselines, RadMamba improves accuracy by 9.4 percentage points while using 86.5 k fewer parameters and the same 6.0 M #FLOP/Inf.
Figure 6: Accuracy vs. number of parameters (1) and number of #\#FLOP/Inf. (2) on: (a) Dataset DIAT (b) Dataset CI4R (c) Dataset UoG20
On Dataset UoG20, which involves FMCW radar in a continuous setting with 6 activities, RadMamba recorded an accuracy of 89.3% with a standard deviation of $\pm$1.38%, utilizing 6.7 thousand parameters and 1.0M #\#FLOP/Inf. per inference. With an even smaller model size and a more complicated task than in Datasets DIAT and CI4R, RadMamba achieves the best accuracy among the models shown in Tab. V and highlights its effectiveness in continuous HAR scenarios with the smallest model size and #\#FLOP/Inf. per inference.
[…]
We also measured peak memory consumption to address concerns about efficiency. The results in Tab. V reveal relatively small differences between models, with peak CUDA allocated memory ranging from 9.7 MB to 22.8 MB. This limited variation, despite RadMamba’s significantly smaller parameter count, highlights a limitation of benchmarking on general-purpose, throughput-oriented embedded GPU architectures for edge inference when the batch size equals 1. On such platforms, the memory footprint is dominated by system-level overheads, such as CUDA context initialization, driver memory pools, and memory management, which mask the model’s intrinsic parameter efficiency.
The true advantage of RadMamba’s parameter efficiency is its suitability for future hardware-software co-design. Its minimal footprint, ranging from 6.7k to 71.4k parameters across the three tested datasets, is intended to enable the entire model to be buffered within on-chip SRAM on a dedicated RadHAR accelerator. This approach would eliminate power-hungry off-chip DDR memory accesses, a primary source of power consumption in on-sensor systems. This fundamental power-saving benefit, along with the low arithmetic utilization for batch-size-1 inference on GPUs, cannot be captured by measurements on the Jetson platform. Therefore, we contend that the #\#params and #\#FLOPs metrics reported in Tab. V are more representative indicators of the RadMamba’s potential for enabling ultra-low-power, on-sensor inference, which is the primary goal of this work.
VI-B Model Scaling and Efficiency Analysis
The model size swept results, illustrated in Figure 6, compare accuracy against the number of parameters from 2k to 150k and #\#FLOP/Inf. across Datasets DIAT, CI4R, and UoG20 between previous models ResNet6, Bi-LSTM, CNN-Bi-GRU, and CNN-LSTM and the proposed RadMamba.
For Dataset DIAT, as presented in Fig. 6 (a1) and (a2), RadMamba achieves a peak accuracy of 99.8% with only 21.7 k parameters. However, RadMamba overfits with a model size larger than 21.7 k because the dataset is very simple and the accuracy is close to 100%. In contrast, ResNet6’s accuracy increases steadily with model size, reaching 99.9% with 208.1 k parameters, yet it requires nearly ten times the parameters of RadMamba for slightly lower performance.
On Dataset CI4R, as shown in Fig. 6 (b1) and (b2), RadMamba consistently outperforms previous models across a range of parameter configurations in efficiency, maintaining superior accuracy with fewer parameters and #\#FLOP/Inf. Similarly, for Dataset UoG20 shown in Fig. 6 (c1) and (c2), RadMamba still records a higher accuracy for continuous activities scenarios with fewer #\#FLOP/Inf., outperforming larger and more computationally intensive models in efficiency and effectiveness.
These trends, as depicted in the figure, affirm RadMamba’s capability to achieve superior accuracy across diverse radar modalities while drastically reducing both parameter count and #\#FLOP/Inf., positioning it as an optimal solution for lightweight, real-time human activity recognition.
VI-C Ablation study of the Micro-doppler-oriented Structure
@@
TABLE VI: Ablation Study with progressive techniques (denoted by Tech.) of New Blocks in RadMamba across Seed 0 to 5, Evaluated with Three Different Radar Settings alongside Their Number of Parameters
Dataset DIAT Dataset CI4R Dataset UoG20
@@
    $(H_{1},W_{1})$ equals to (7, 7) for Dataset DIAT and B, (5, 7) for Dataset UoG20 due to the different image size of these datasets.
@@
    The kernel size for projection 1 and 2 is $(dim,dim,3)$, and for projection 3 is $(dim,dim,1)$.
@@
Table VI presents an ablation study evaluating the impact of new architectural blocks in the RadMamba model across three datasets with varying radar settings. The study systematically assesses the effects of different projection layers, patch sizes, and downsampling scales on model accuracy and parameter efficiency, with results averaged over seeds 0 to 5.
@@
The main conclusion from this ablation study is that optimal performance emerges from the combined use of all three proposed new blocks, including the projection layer, the patch size, and the downsampling scale, rather than any single component in isolation. This is evident in the incremental improvements observed across the configurations. For instance, the baseline conventional ViM model (Row 1) achieves an accuracy of 93.5% on Dataset DIAT, 61.9% on Dataset CI4R, and 71.2% on Dataset UoG20. In contrast, the proposed RadMamba (Row 27), which incorporates a 1D convolution layer, a patch size of ($H_{c-d}$, 1), and a downsampling (2, 2), achieves a higher accuracy of 99.8% on Dataset DIAT. Similarly, on Dataset CI4R, Row 27 with a (2, 8) downsampling scale yields significantly higher accuracy of 89.6%, and on Dataset UoG20 with a (2, 32) scale, it reaches 88.6%. Intermediate configurations, such as Row 19 (only projection replacement), Row 7 (only patch size changing), and Row 2 (only downsampling applied) show lower accuracies—underscoring that no single structure alone achieves the peak performance of the fully combined design in Row 27.
@@
To provide a more detailed analysis, we discuss the three blocks—projection layer, patch size, and downsampling scale—separately, evaluating their individual contributions to RadMamba’s performance across the datasets.
VI-C1 Projection Replacement